OpenStudy (anonymous):

Given the function g(x) = 8x - 2, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values. The value of g(-2) is larger than the value of g(4). The value of g(-2) is the same as the value of g(4). The value of g(-2) is smaller than the value of g(4). The values of g(-2) and g(4) cannot be compared.

OpenStudy (ipwnbunnies):

Plug in -2 for x. Plug in 4 for x. Compare the y-values you get for both inputs.

OpenStudy (ipwnbunnies):

g(-2) = 8(-2) - 2 g(4) = 8(4) - 2

OpenStudy (anonymous):

so A

OpenStudy (ipwnbunnies):

Err, no...What's g(-2) and what's g(4)...

OpenStudy (ipwnbunnies):

I plugged in the x-values for you. You do the math.

OpenStudy (anonymous):

Okay now im thinking C

OpenStudy (ipwnbunnies):

Why are you thinking C?

OpenStudy (anonymous):

because with the equations you gave me it seems to make sense idk im probably wrong

OpenStudy (ipwnbunnies):

What? Just. Do. The. Math. g(-2) = 8(-2) - 2 = ?? What does that equal?

OpenStudy (anonymous):

The value of g(-2) is smaller than the value of g(4) With g(-2) you get -18 With g(4) you get 30

OpenStudy (anonymous):

So I was right it's C!

OpenStudy (anonymous):

Thank you thank you

OpenStudy (ipwnbunnies):

You were right you guessed.

OpenStudy (anonymous):

It wasn't a guess and you don't need to be rude